A boat goes 15 miles per hour in still water. If it can go 12 miles downstream in the same amount of time as it takes to go 9 miles upstream, then what is the speed of the current?
Answer by stanbon(75887)

A boat goes 15 miles per hour in still water. If it can go 12 miles downstream in the same amount of time as it takes to go 9 miles upstream, then what is the speed of the current?
-----------------------------
Down-stream DATA:
distance = 12 miles ; rate c + 15 mph ; time = 12/(c+15) hrs.
-------------------------------
Up-stream DATA:
distance 9 miles ; rate = 15-c mph ; time = 9/(15-c) hrs.
-------------------------------
Equation:
time = time
12/(15+c) = 9/(15-c)
12(15-c) = 9(15+c)
12*15 - 12c = 9*15 + 9c
21c = 3*15
c = 15/7 = 2 1/7 mph (speed of the current)
===============================================
Cheers,
Stan H.
=============